Find an equation of the line passing through the given points. Use function notation to write the equation.

​(3​,11​) and ​(4​,16​)


1
Expert's answer
2022-02-14T16:37:43-0500

Let's compose the equation of the line at two points:

"\frac{{x - 3}}{{4 - 3}} = \frac{{y - 11}}{{16 - 11}} \Rightarrow 5(x - 3) = y - 11 \Rightarrow y = 5x - 15 + 11 \Rightarrow y = f(x) = 5x - 4"

Answer: "f(x) = 5x - 4"